OpenStudy (anonymous):

Find the mean and the standard deviation to the nearest hundredth for each set of data. Justify your reasoning. {302, 310, 331, 298, 348, 305, 314, 284, 321, 337}

OpenStudy (amistre64):

those number are too big, id subtract say, 305 from all of them

OpenStudy (amistre64):

{302, 310, 331, 298, 348, 305, 314, 284, 321, 337} - 305 {-3, 5, 26, -7, 43, 0, 9, -21, 16, 32} that should be more doable

OpenStudy (amistre64):

add em up ....

OpenStudy (anonymous):

for the mean I just add them and divide by 10? but how do I find the standard deviation to the nearest hundredth? @amistre64

OpenStudy (amistre64):

we would need the mean in order to do that, so lets get the mean of the smaller set

OpenStudy (amistre64):

sum of {-3, 5, 26, -7, 43, 0, 9, -21, 16, 32} = 100 ; mean = 100/10 = 10 recall that we subtracted 305 from all the values, so we would add 305 back to it; giving us a mean of 10+305 = 315

OpenStudy (amistre64):

the standard deviation of the smaller set is the same as the larger set; they have the same relative spread .... to find the standard deviation of a set, we subtract the mean from each value to begin with ... notice that the smaller set is already -305 and we would just have to -10 more - which is the mean of the smaller set to begin with

OpenStudy (anonymous):

oh okayyy yeahh. I got the mean but the only thing I don't understand is the standard deviation part

OpenStudy (amistre64):

since we are already -305 on the smaller set, lets continue with the -10 (which is the mean of the smaller set to begin with) {-3, 5, 26, -7, 43, 0, 9, -21, 16, 32} - 10 {-13, -5, 16, -17, 33, -10, -1, -31, 6, 22}

OpenStudy (amistre64):

now we just need to square these results ... and get the mean of the squares

OpenStudy (amistre64):

{-13, -5, 16, -17, 33, -10, -1, -31, 6, 22} ^2 {169, 25, 256, 289, 1089, 100, 1, 961, 36, 484}

OpenStudy (amistre64):

again we could subtact say 150 seems reasonable {169, 25, 256, 289, 1089, 100, 1, 961, 36, 484} - 150 {19, -125, 156, 139, 939, 0, -149, 811, -114, 334} {2010}/10 = 210 210+150 = 360; sqrt(360) = 6 sqrt(10) but since i did that by head .... let me dbl chk

OpenStudy (amistre64):

yeah, i added wring :) 3410/10 = 341, so sqrt(341) is best

OpenStudy (anonymous):

thank you so much that helped me out a lot! @amister64